Calculate the speed of sound in air at 42°C.
Calculate the speed of sound in seawater at 20°C given that the bulk modulus of seawater is 2.34×10^9 N/m² and its density is 1.025×10³ kg/m³.
Calculate the speed of an aircraft with a mach number of 1.23. Assume that the speed of sound is 340 m/s.
1
Expert's answer
2018-11-12T15:57:09-0500

c= 331.3 + (0.6)T= 331.3 + (0.6)42= 356.5 m/s.


Calculate the speed of sound in seawater at 20°C given that the bulk modulus of seawater is 2.34×10^9 N/m² and its density is 1.025×10³ kg/m³.

Solution

c=√((2.34∙〖10〗^9)/(1.025∙〖10〗^3 ))=1510 m/s.


Calculate the speed of an aircraft with a mach number of 1.23. Assume that the speed of sound is 340 m/s.

Solution

v=1.23c=1.23(340)=418 m/s.